OpenStudy (anonymous):

mx'' = gm - kx' Come up with a change of variables to solve this IVP, giving an equation for the position of the object at time t.

OpenStudy (anonymous):

try V= dx/dt and dV/dt = d^2x/(dt^2)

OpenStudy (anonymous):

hmm ... ok

OpenStudy (anonymous):

\[\frac{ dV }{dt }+ \frac{ k }{m }V =g\] \[\huge u = e ^{\int\limits_{ }^{ }\frac{ k }{m }dt}\] \[\huge u =Ce ^{\frac{ kt }{m } }\] \[ Ce ^{\frac{ kt }{m } } *(\frac{ dV }{ dt } +\frac{ k }{m } V =g)\]

OpenStudy (anonymous):

seen this method before? integrating factor?

OpenStudy (anonymous):

yeah i have

OpenStudy (anonymous):

k, got it from here?

OpenStudy (anonymous):

think so, thank you

OpenStudy (anonymous):

@Algebraic! i dont have this as much as i thought i did!

OpenStudy (anonymous):

did you find V?

OpenStudy (anonymous):

from the step where we multiplied through by u:

OpenStudy (anonymous):

yeah, ive done that.

OpenStudy (anonymous):

\[Ce ^{\frac{ kt }{ m }} \frac{ dV }{ dt } + Ce ^{\frac{ kt }{ m }} \frac{ k }{ m }V = Ce ^{\frac{ kt }{ m }} g\]

OpenStudy (anonymous):

how far have you gotten?

OpenStudy (anonymous):

ya i have that

OpenStudy (anonymous):

LHS now looks like the result of an application of the product rule... ie V' *u +V*u'

OpenStudy (anonymous):

rewrite as (V*u)' = u*g

OpenStudy (anonymous):

integrate both sides

OpenStudy (anonymous):

\[ Ce ^{\frac{ kt }{ m }}V =\int\limits_{ }^{ } Ce ^{\frac{ kt }{ m }} g\]

OpenStudy (anonymous):

ok wait ..

OpenStudy (anonymous):

why do we insert u*g ?

OpenStudy (anonymous):

we found u so that the LHS would look like the result of a product rule : (V*u)' once we multiplied the *whole* equation by u

OpenStudy (anonymous):

so I didn't 'insert' ug , we multiplied everything by u...

OpenStudy (anonymous):

uV' + k/m *u*V = gu

OpenStudy (anonymous):

right ok

OpenStudy (anonymous):

ok, so how did you know (vu)' = gu ?

OpenStudy (anonymous):

i think i'm slow . idk why i'm taking so long to get this

OpenStudy (anonymous):

when we integrate both sides for the last part its by dt right?

OpenStudy (anonymous):

ok.. so i've made it to \[Ce^(kt/m) * v = mg/k * Ce^(kt/m) \]

OpenStudy (anonymous):

@amistre64 @Hero could you possibly help me with where i should go with this next step?

OpenStudy (anonymous):

right so..sorry I got called away abruptly.

OpenStudy (anonymous):

thats ok, thanks for coming back

OpenStudy (anonymous):

but you're integrating from t=0 to t=t, so you get\[\huge Ce ^{\frac{ kt }{ m }} V = \frac{ mg }{ k }(Ce ^{\frac{ kt}{ m }} -Ce ^{\frac{ k0}{ m }} )\]

OpenStudy (anonymous):

divide through by C*e^(kt/m) and you've found V

OpenStudy (anonymous):

integrate V(t) to get x(t)

OpenStudy (anonymous):

ok .. trying that

OpenStudy (anonymous):

if you ever come back, i'd love your help

OpenStudy (anonymous):

I'm here, what's up?

OpenStudy (anonymous):

AH, sorry, i missed you!! Ok, so i've done everything you said. Ce^kt(0)m is just a constant when you integrate it right?

OpenStudy (anonymous):

well...

OpenStudy (anonymous):

did you divide through by Ce^(kt/m) yet?

OpenStudy (anonymous):

ya .. so it isn't when you go to integrate it because its still over the other equation

OpenStudy (anonymous):

go up to the last post I made with an equation. that's after integrating. divide both sides by Ce^(kt/m) now you have V

OpenStudy (anonymous):

get V and then we'll go on and get x(t)

OpenStudy (anonymous):

wait .. we have to find v, then x(t) ? i thought by dividing through we found v, then we had to integrate that .

OpenStudy (anonymous):

to get x(t)

OpenStudy (anonymous):

you do, but you're skipping steps or something

OpenStudy (anonymous):

you're evidently not getting the right V(t). tell me what you get for V(t)

OpenStudy (anonymous):

hmm .. well if you just divide (Ce^kt/m) through i got .. v = mg (Ce^(kt/m) - Ce^(kt(0)/m)) // kCe^(kt/m)

OpenStudy (anonymous):

then we're supposed to integrate that, right?

OpenStudy (anonymous):

simplify it

OpenStudy (anonymous):

Ce^0 = C C/(Ce^(kt/m) ) = e^(-kt/m)

OpenStudy (anonymous):

(Ce^(kt/m) )/(Ce^(kt/m) ) =1

OpenStudy (anonymous):

\[\huge V = \frac{ mg }{ k } (1-e ^{\frac{-kt }{m} })\]

OpenStudy (anonymous):

why is it -k?

OpenStudy (anonymous):

because it's more fun.

OpenStudy (anonymous):

than k

OpenStudy (anonymous):

ok, so now we integrate that?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

wait, can i ask why its 1 - e^(kt/m)

OpenStudy (anonymous):

you could ask that, but it's not 1 - e^(kt/m) so it would be fruitless...

OpenStudy (anonymous):

hmm ok... i'm not sure if i integrated properly. but i did it on the interval 0 - t again because of the second initial condition ... right? and got mgt/k - (m^2g/k^2)(e^kt/m) - m^2g/k^2

OpenStudy (anonymous):

naw...

OpenStudy (anonymous):

what are you integrating?

OpenStudy (anonymous):

\[\huge \frac{ mg }{ k } \int\limits_{0}^{t} (1-e ^{\frac{ -kt }{m }}) dt\]

OpenStudy (anonymous):

yeah. thats what i integrated, but i left k postiivive.. can i do that?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

oh so it has to be ive? its not just for the fun of it?

OpenStudy (anonymous):

ngative i meant aha.

OpenStudy (anonymous):

C/(Ce^(kt/m) ) = e^(-kt/m)

OpenStudy (anonymous):

so that happens for anything you divide into the numerator then?

OpenStudy (anonymous):

\[\huge \frac{ 1 }{ x ^{a} } =x ^{-a}\]

OpenStudy (anonymous):

idk why this is so tough for me, thanks for being so patient. ok, so i'm integrating that now.

OpenStudy (anonymous):

cool:)

OpenStudy (anonymous):

so i would get t + m/ke^(-kt/m)

OpenStudy (anonymous):

right?! then evaluate it?

OpenStudy (anonymous):

SO AS AN ANSWER we would get x(t) = t - (m/k)(e^(-kt/m)) - m/k

OpenStudy (anonymous):

RIGHT?!

OpenStudy (anonymous):

please tell me i`m right. i dont know what i'll do if i'm not

OpenStudy (anonymous):

ya pretty close (mg/k) (t + (m/k)(e^(-kt/m)) - m/k) if that sign in front of the exponential isn't positive, then at t=0 you're not at x=0

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

oh ya thats what i actually had in my notes .. thank goodness. thank you so much for everything!

OpenStudy (anonymous):

Sorry, one more question. In this, m = mass, gravity = g and k = acceleration right?

OpenStudy (anonymous):

so if its asking what the terminal velocity is of the falling object, i just make g = 9.8 and k = 0 right?

OpenStudy (anonymous):

I am not getting notifications from you... the server must be stuffed.

OpenStudy (anonymous):

you'll have to PM me if you want attention..

OpenStudy (anonymous):

yeah m = mass g= acceleration k is just a constant relating to aerodynamics, cross section and shape etc

OpenStudy (anonymous):

so would i evaluate x, or find it's first der. then evaluate with those constantS?

OpenStudy (anonymous):

you have it all now, x(t), V(t) you can find a(t) easily (derivative of V(t) )... you can do whatever you want....!

OpenStudy (anonymous):

you already have V(t) so don't bother taking a derivative of x(t) unless you're a masochist.

OpenStudy (anonymous):

i try not to be

OpenStudy (anonymous):

so plug them into v then

OpenStudy (anonymous):

w. k being 0, it's undefined .. right? .. that doens't seem right

OpenStudy (anonymous):

k=0?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

for acceleration? no?

OpenStudy (anonymous):

or is k resistance?

OpenStudy (anonymous):

if that's true the whole problem was pretty pointless...